1) A car is initially moving at 35 m/s is accelerated at 4.0 m/s2 for a period of 7s. The car then travels with the speed after 7s for another 10s. What is the average speed of the car for the 17 s period?

a.    63 m/s

b.    57 m/s

c.    49 m/s

d.    35 m/s
Answer #1

V = u +at = 35 + (4 x 7) = 63 m/s

average velocity = total dislacement / total time

for 1st 7 sec

S = ut +1/2 at^2 = (35 x 7) + (1/2 x 4 x 7^2) = 343 m

for 10 sec

S = ut + 1/2 at^2 = ut + 0 = 63 x 10 = 630 m

Now

average velocity = (630 + 343) / 17 = 57 m/s (option b)